Project

General

Profile

Task #22808

5GHz, Cavity, (TE5RI01) + quantum amplifiers

Added by Daniil Frolov 8 months ago. Updated 7 months ago.

Status:
New
Priority:
Normal
Assignee:
-
Start date:
06/24/2019
Due date:
% Done:

0%

Estimated time:
Duration:

6272019618_switch.JPG (81.3 KB) 6272019618_switch.JPG Daniil Frolov, 06/27/2019 06:18 PM
6272019642_switch.JPG (80.1 KB) 6272019642_switch.JPG Daniil Frolov, 06/27/2019 06:44 PM
62820191023.PNG (39.1 KB) 62820191023.PNG Daniil Frolov, 06/28/2019 10:27 AM
field_gui_2019_Jun_28__16h17m25s.jpg (211 KB) field_gui_2019_Jun_28__16h17m25s.jpg Daniil Frolov, 06/28/2019 06:33 PM
6282019635.JPG (120 KB) 6282019635.JPG Daniil Frolov, 06/28/2019 06:36 PM
amp_on160407012019.png (66.2 KB) amp_on160407012019.png Daniil Frolov, 07/01/2019 04:02 PM
twpa_disp_feature_712019736.png (56.5 KB) twpa_disp_feature_712019736.png Daniil Frolov, 07/01/2019 07:41 PM
dual_gui_2019_Jul_01__19h35m06s.jpg (392 KB) dual_gui_2019_Jul_01__19h35m06s.jpg Daniil Frolov, 07/01/2019 07:47 PM
dual_gui_2019_Jul_01__20h13m04s.jpg (239 KB) dual_gui_2019_Jul_01__20h13m04s.jpg Daniil Frolov, 07/01/2019 08:22 PM
twpa_784_gain_712019928.png (274 KB) twpa_784_gain_712019928.png Daniil Frolov, 07/01/2019 09:29 PM
TWPAstudy752019508.png (203 KB) TWPAstudy752019508.png Daniil Frolov, 07/05/2019 05:19 PM
752019826_1.png (95.2 KB) 752019826_1.png Daniil Frolov, 07/05/2019 08:34 PM
071120191055.png (63.7 KB) 071120191055.png Daniil Frolov, 07/11/2019 10:50 AM
07112019111109.png (71.8 KB) 07112019111109.png Daniil Frolov, 07/11/2019 11:05 AM
71120191157.png (72.8 KB) 71120191157.png Daniil Frolov, 07/11/2019 11:57 AM
11527252019.JPG (124 KB) 11527252019.JPG Daniil Frolov, 07/25/2019 11:52 AM

History

#1 Updated by Daniil Frolov 8 months ago

Found resonance frequency 5.044386 GHz @ 3.75 K. Qis ~7E+7 from 3dB bandwidth measurement.

#2 Updated by Daniil Frolov 8 months ago

Started LabView program for Q vs temperature measurement at ~5 pm. At 7-30 pm I found that spectrum analyzer software crashed. I restarted LabView program and spectrum analyzer.

#3 Updated by Daniil Frolov 8 months ago

  • Subject changed from Measure Q vs temperature (5GHz, TE5RI01) to 5GHz, Cavity, (TE5RI01) + quantum amplifiers

#4 Updated by Daniil Frolov 8 months ago

I have just turned on control of switch #1, that is located right after the 5 GHz cavity. I started control app and tried to read the state of the switch. It reads state correctly, and doesn't heat up. At least 10 minutes after readout command, temperature didn't change at all. A will now try to activate the switch, and turn it to different position...

#5 Updated by Daniil Frolov 8 months ago

...I activated the switch, it works!!! As we expected, temperature of mixing chamber jumped up to ~90 mK. It took ~25 minutes to cool back to 20 mK.

#6 Updated by Daniil Frolov 8 months ago

Some update on switching. Yesterday I forgot to use current-limiting resistor in series with the switch coil, that we planned to use when the wire of the coil will become superconducting, so the current through the coil was very high.

I tried to activate switches with current-limiting resistor this morning, and it works! Now the temperature jumps very little. Below is temperature of mixing chamber after two switches were activated, one after another.

First, temperature jumped from 12.5 to 15.5 mK, and next time it jumped from 15.5 to 18.5 mK - very small change.

Looks like that switches heat mixing chamber negligibly small.

#7 Updated by Daniil Frolov 8 months ago

I'm starting to tune Yale quantum amplifier. Finished flux scan with new program. Found optimal biasing at +230 uA. Quantum amplifier resonates at 4.984 GHz with this biasing.

#8 Updated by Daniil Frolov 8 months ago

Finally got some gain from the Yale amplifier (~30 dB). Pump settings are 13.940 GHz, +1.9 dBm. Bias is the same, +230 uA.

#9 Updated by Daniil Frolov 8 months ago

I started to activate switches and temperature went up to 36 mK, then the gain from the quantum amplifier disappeared. I re-scanned it and found that the magnetic field vs. amplifier resonant frequency spectrogram also shifted. I foud the new parameters of the amplifier:
biasing +264 uA
pump: 14.160 GHz
pump level: +0.5 dBm

with these parameters Yale quantum amplifier gives 35 dB gain at 4.9 GHz and -70 dBm input power (to the port 8 of the fridge).

#10 Updated by Daniil Frolov 8 months ago

Started working with TWPA. #784.
Found dispersive feature according to manual.

Made a fast spectrogram, pump frequency (qubit vertical axis in the picture) vs. TWPA transfer function ("Cavity" horizontal axis in the picture). Looks like the desired, wide gain profile is achieved when pump frequency is around 6300 MHz (1 dbm power setting on the generator).

Now will make a finer scan, to measure pump frequency more accurately.

#11 Updated by Daniil Frolov 8 months ago

Finished fine scan. From the scan best pump frequency is 6.287 GHz.

#12 Updated by Daniil Frolov 8 months ago

Finally measured TWPA gain with various levels of pump power.

#13 Updated by Daniil Frolov 8 months ago

Summary on TWPA # 784:

It works, however behavior with the cavity is different from the behavior when it is connected to the source directly, without cavity. When the amplifier is tuned without cavity gain can reach 20 dB @ 5 GHz with the specific settings of pump frequency and power. However, when the cavity is connected to the input of the amplifier, gain with the same settings of pump signal disappears. With the cavity connected the amplifier requires different setting of pump frequency and power, so it moves to some different operating point and in this mode maximum gain is not achieved any more.

In the original setting without cavity, with settings of the pump= 6.287 GHz and +0.4dBm gain was around 20 dB at 5 GHz. When I connected cavity, I found maximum gain 8 dB was at pump = 6.274 GHz, - 4 dBm. So with cavity gain of the TWPA drops by at least 10 dB. This may be because of the mismatch between cavity output and TWPA input: we don't have any isolators there currently.

Also I was able to observe ~8 dB improvement in SNR with TWPA for the resonance peak measurement (see upper plot).

#14 Updated by Daniil Frolov 8 months ago

  • File 752019826.png added

#15 Updated by Daniil Frolov 8 months ago

Just finished measurements with the second TWPA # 600. Results are very similar. Pump frequency and power is slightly different: 6.317 GHz, -2 dBm for optimal gain around 5 GHz.

The same problem with low gain when connected to the cavity remains. Next run need to add idolators between cavity and TWPA.
Also see improvement in SNR when TWPA is on, about 8 dB improvement

#16 Updated by Daniil Frolov 8 months ago

  • File deleted (752019826.png)

#17 Updated by Daniil Frolov 7 months ago

Haozhi is here. We are now starting to test NIST JPA.

We made manual scan with network analyzer and found pump frequency (bifurcation point)= 5.04 GHz at bias -7 uA, -30.8 dBm power setting on the VNA (9.2 + 40 external attenuator).

#18 Updated by Daniil Frolov 7 months ago

Turned on pump...
See gain = ~25 dBm.
VNA output power is -70 dBm (-30dBm + 40 dB attenuation).
Pump power = -29.44 dBm (Rohde&Schwarz generator), pump frequency= 5.04 GHz.

#19 Updated by Daniil Frolov 7 months ago

Connected cavity 5GHz to the amplifier.
Slightly changed pump parameters to optimize gain. 5.0446 GHz, power= -29.34 dBm (R&S generator).

observing gain about 25 dB!

On picture: saved trace (red)- pump off, active trace (yellow)- pump on.

#20 Updated by Daniil Frolov 7 months ago

Update:
we tried to measure noise temperature of the system with JPA and total gain of the system.
We used y-factor method with calculated quantum noise as cold source and matched load sitting at 1.002 K as a hot source (connected to JPA with superconducting wire through the switch). We studied next configurations:

1) JPA was off, its input connected to the matched load at physical temperature of 20 mK. Half-photon quantum noise temperature is 0.24 K @ 5.004 GHz. In this case total system noise should be dominated by thermal noise of the HEMT (2.6182 K). Measured output power at the spectrum analyzer (JPA off + HEMT + room temperature amplifier) was -129.46 dBm (zero span, 1 Hz bandwidth).

2) JPA was on, under the above conditions output power on the spectrum analyzer increased to -116.16 dBm.

So the difference in output power between the two conditions JPA "on" and JPA "off" is -129.46 dBm -(-116.16 dBm)= -13.3 dBm
Theoretical difference between HEMT's temperature noise power and half-photon quantum noise power is -207.145 dBm - (-194.619 dBm) = -12.52 dBm, which is only 0.78 dB off the difference that we measure.

Our estimation of total gain of the whole chain from JPA input to the spectrum analyzer is 91 dB based on the measurements above.
(See python file with calculations in today's shift folder on RF computer: c:\redmine\runs\DR1_run_5\python_measurements\2019_07_11\NIST_JPA\JPA_noise_studies

3) We tried to measure effective noise temperature with y-factor method, by connecting "hot" load @ 1.002 K to the JPA. Output power measured with the spectrum analyzer in this case was -115.06 dBm. Effective noise temperature calculated with y-factor method was 2.39 K, which is very high and seems to be wrong (should be order of 20 mK ?). To reach 20 mK one needs to increase the result of output power with the hot load by factor of ~3 which is 4.7 dB. This means that we are probably loosing 4.7 dB on the way from hot load to the input of JPA, which seems relatively high. Need to study this more.

#21 Updated by Daniil Frolov 7 months ago

Some results from low power cavity decay measurements.

Yesterday we tuned JPA and measured several decays with and without averages. Attached chart shows estimation for photon levels at the output of the cavity, based on the assumption that when no signal is applied to the input of the cavity, JPA amplifies only half-photon noise. Decay curve and noise curve on the right is measured with 20 averages.
The gain of the chain is 77.5 dB and was estimated as a ratio between the noise level measured on the spectrum analyzer when JPA was on and calculated half-photon noise. Noise of the JPA is not taken into account in this estimation.
Mixing chamber temperature was 23 mK.
Data is stored here: \\tdserver1\project\Quantum Computing\redmine\runs\DR1_run_5\manual_measurements\TE5RI001_decay_with_JPA



Also available in: Atom PDF